# The Atacama Cosmology Telescope: The Development of Machine Learning Tools for Detecting Millimeter Sources in Timestream Pre-processing

### v1

@@ 4.2.2 Normalized Maximum
To derive this statistic, we compute the histogram of the distribution of the detectors in the direction of the focal plane for a given glitch. The histogram bin with the largest number of detectors and the adjacent bins are added together and then normalized by the total number of detectors present in the glitch. We compute the histogram of detectors with 10 bins, resulting in the width of an individual bin in the histogram being small when the detectors are clustered and larger if there is spread in positions. Figure 3 shows examples for both the PS and PS+ categories, where the detectors being considered are shown as coral compared to all the detectors in the snippet shown in purple. Note that in both examples, there is only one bin beside the maximum bin, thus only two bins are used for the “Normalized Maximum” statistic. For the PS+, this is because the maximum bin is the first bin, whereas for the PS it is due to the detectors that would have been in the bin to the right of the maximum being cut during the detectors cuts which happen before the glitch cuts. The PS object, has only 27% of the detectors within the maximum and adjacent bins, compared to 59% of detectors within the maximum and adjacent bins for the PS+ object.
@@
In general, we expect the spread of affected detectors in the direction to fall within 0.1, or a few beams, for an astrophysical source. This is not true for a general electronic glitch or cosmic ray, which makes this metric useful for identifying astrophysical sources. We compute the fraction of detectors in the direction of the focal plane that are within 0.1 of the histogram bin that contains the maximum number of detectors. Figure 3 shows examples for both the PS and PS+ categories for both the “Normalized Maximum” and “Normalized Maximum Within 0.1” to compare these statics.
Similarly to the previously defined statistic, we expect a small range of detectors to observe a point source as the telescope scans over the source. Therefore, the number of detectors within 0.1 of the position of the maximum histogram bin should be close to the total number of detectors in the snippet (and therefore our statistic normalized over the total number of detectors should be very close to unity, as shown in Figures 2 and 3). This makes the statistic very useful for determining if a glitch is a point source. A cosmic ray will hit a small cluster of detectors, resulting in a small spread in the direction. We still expect the fraction of the total number of detectors within 0.1 of the position to be close to unity, but there can be deviations from this if multiple cosmic rays hit the focal plane at once, and they are still generally smaller than for a point source. The expected distribution of detectors varies quite significantly depending on the type of electronic glitch is occurring, but we find that the statistic is generally smaller for electronic glitches than for the PS class. For a PS+ object, the value of this statistic will differ from what is seen for an isolated source depending once again on where the additional glitch occurs on the focal plane relative to the source location: if the glitch occurs near the source, we similarly expect many detectors to be included, leading to a large fraction of the detectors to be within 0.1 of the maximum. Conversely, if the glitch occurs far away on the focal plane, this will result in a smaller fraction of detectors within 0.1 of the maximum and a smaller value for our statistic. This statistical metric is therefore useful in identifying PS+ objects where the coincident glitches are located elsewhere in the detector array and thus would be able to be easily separated in the snippets from the point source and for further analysis of the source.
@@ 4.3 Correlation
In order to determine how the incident flux is correlated across detectors (which can help distinguish between the PS and CR classes), we compute the Pearson correlation coefficient of each detector pair ordered by their position across the focal plane. We then use the mean of the absolute value of these correlation coefficients from all detector pairs in the snippet as a feature in our classification algorithm. Given that detectors observe the point source sequentially as the telescope scans across the sky, we only expect a large value of this correlation coefficient between detectors that are either close to or beside the main detectors on the focal plane that observe the source. This scenario leads to a block diagonal correlation matrix of high values across the array when the detectors are ordered by their position (see Figure 4 for an illustration) but a low overall degree of correlation across all detectors. For cosmic rays, we expect a high level of correlation across the array, because the detectors are all affected by the cosmic ray at the same time, shown in Figure 4. For electronic glitches that cause a strong signal that affects multiple detectors simultaneously, we expect a large value of the correlation coefficient during, e.g., a large-area heat-up. Conversely, if the glitch is similar to a random noise fluctuation, we would expect a low value of the correlation coefficient. For PS+ objects, if the additional glitch is a cosmic ray, we expect a slightly higher correlation than would be observed for a point source, as the cosmic ray portion of the TODs for each detector will overlap while affected by the cosmic ray. If the additional glitch is of electronic or other origin, the effect varies, but it generally still increases the level of correlation, as shown in Figure 2.
@@ 4.4 Time-lag
While the absolute value of the correlation described above indicates the spatial correlation between detectors that see a given glitch, the temporal correlation (or time-lag) is also a key statistic that we use in classification. The time-lag between detectors is the shift in time that maximizes the cross-correlation between the two detector TODs over the snippet. For two detectors and , it is defined following 35,
where DFT is the discrete Fourier transform, the overbar denotes the complex conjugate, and is the timestream of the detector as a function of time. We take the absolute value, since the sign of the shift (backwards or forwards) does not contain information relevant to the type of the signal. We compute this time-lag for all detector pairs in the snippet, and then use the mean of the absolute values to summarize the glitch.
For a point source, each detector observes the source sequentially as the telescope scans across the sky. We therefore expect a time-lag that is equal to the separation in azimuth of the two detectors divided by the scanning speed. In practice, we can get different values of the time-lag based on how noisy the TOD is, especially for low amplitude sources. This quantitative number can be used in future work to differentiate an astrophysical point source from a moving source, such as a satellite. Conversely, we expect a time-lag close to zero for cosmic rays, since the cosmic ray timescale is fast compared to the readout rate, which means that the digitized signatures reside essentially on top of each other in the detector timestream. This can be seen in Figure 5, where the time-lag for detector pairs, ordered by their position, are shown for a cosmic ray on the top and a point source on the bottom. As expected, the time-lag is zero for almost every detector pair for the cosmic ray. For the EG category, it will cause varied time-lags depending on the nature of the glitch. Generally, something such as a focal plane heating event or another electronic glitch that affects all the detectors at the same time will result in a zero time-lag as the detectors experience the effect at the same time. Finally, for the PS+ class, if the other glitch is a cosmic ray, we expect a slightly smaller time-lag than just a point source as part of the TODs will be overlapping with a zero time-lag. If the other glitch is of an electronic or other origin the effect varies, but generally speaking it increases the time-lag, as we show in Figure 2.
@@ 4.5 Number of Peaks
In a given snippet, some glitches, in particular glitches from PS objects and CR objects, will appear as “peaks” in the TOD, and so the number of signal peaks at unique times is used as one of our summary statistics, as it is particularly useful in discerning a cosmic ray from an electronic glitch. In order to compute the number of peaks in a snippet, we first smooth the TODs for the snippet using a boxcar smoothing kernel of width three time samples to remove small fluctuations due to random noise. We then combine the TODs from each detector into a single timestream , as:
@@
4.6 Importance of Statistics
When using the various statistics described above, it is important to determine which of the features are more instructive in the classification step, to ensure that we understand how the classifier weights them. Low-importance features can potentially be removed from a feature set, and high-importance features can be examined to identify the physics behind why the features lead to more accurate classifications. It also aids in our understanding of misclassified glitches; see Section 7.1 for more details.
4.7 Optimizing the Machine Learning Algorithm
We explored the performance of the random forest as a function of the number of trees and the “max depth” (which is the maximum number of times each decision tree is allowed to split) of the forest classifier. The number of decision trees used was varied between 25 and 175 (preliminary tests outside that range showed minimal impact in performance) and the max depth varied within an initial range of 2 to 32. The performance was assessed using the overall accuracy of the classifier, the overall out-of-bag (OOB) score, and F-1 score which is the harmonic mean of the precision and recall (see Section 7.1 for more information on our metrics) for the four glitch categories. The random forest approach uses bootstrap aggregation to train each tree, a process whereby the training set for each decision tree is produced by sampling from the original training set with replacement. This avoids overfitting of the algorithm to the training data, which can lead to a model that is does not perform well on new (test) data. The ability of a method like random forests to make predictions on data that were not included in the training set for each individual decision tree within the ensemble of trees is characterized by the OOB score 24. The performance metrics discussed above plateau and do not improve when the forest reaches a size of 50 trees and a max depth value of 15; therefore, we fixed those hyperparameters of the random forest method to these values.
5 Training and Test Set Preparation
@@ 5.1 Labeling of Objects in the Training and Test Sets
Conventionally, the creation of training and test data proceeds via human (manual) labeling. In order to create a large enough training set for the random forest classifier, we needed an efficient method to facilitate manual labeling of the snippets, via a graphical user interface (GUI).
We use the Python-driven platform developed by the Zooniverse collaboration111www.zooniverse.org to create the labeling GUI. This Zooniverse platform includes a well-established data pipeline and a user-friendly interface, as can be seen in Figure 8. The platform we employ was designed specifically for use by members of the ACT collaboration who can simultaneously label glitches stored on a common computing facility, to assist in both training and verification of the algorithm.
To facilitate classifications by non-expert users, we created a user tutorial specific to ACT glitch data, with examples related to the key features of each type of glitch and notes on pitfalls in classifying glitches.222We separately included instructions on setting up a Zooniverse project from scratch, see online documentation here.
@@ 5.2 Forming the Training and Test Sets
Once we have obtained a representative sample for training and testing our classifier, we need to ensure we are robust to changes in the snippets from different years (observation seasons), observational frequency bands, detector array configurations, etc., which requires further refinement of our test/training split. In order to ensure we are robust to these changes, we use an active learning prescription, in particular the modAL 18 learning framework. Active learning is a machine learning approach where unlabeled objects are selected for manual classification/labeling depending on specific criteria of the object (for example the classification ambiguity between types in a given set, or an estimate of the Bayesian classification of the object type), as opposed to selecting the point for labeling randomly from the unlabeled set. We employed the modAL package to choose objects using a selection criterion based on which objects are most ambiguous, or in other words, objects for which the maximum classification probabilities are the lowest. In doing so, we can leverage human/manual classifications where our random forest classifier might be most uncertain. This criterion for selecting objects for manual classification is known as “uncertainty sampling”. Figure 9 shows a diagram of the process.
@@
From the initial random forest classification, each glitch has four probabilities related to the four possible object types: PS, PS+, CR, and EG. The most likely (highest) probabilities for all glitches are ranked, and the uncertainty sampler takes the lowest value of those probabilities. In other words, the uncertainty sampler selects a glitch that the random forest classifier is the least confident in (or most uncertain about) to manually label and therefore add to the active learning training set 30. This is shown in the bottom right-hand section of Figure 9.
@@
To demonstrate the effects of active learning, we tested the algorithm on a set of 851 labeled glitches from 2021 at 90 GHz and 150 GHz for detector arrays PA4 and PA5 (of which 63 are PS, 44 PS+, 363 CR, and 381 EG). We begin with a training set composed of 2280 glitches from 2019 at 90 GHz in PA5. We added 300 training points using two methods, active learning and then random selection, that were chosen from a set of 2555 glitches from 2021 at 90 GHz and 150 GHz in PA4 and PA5. It is important to note that the test set and additional training points are comprised of data from different arrays, frequencies, and years which could have different noise properties or detector performance than the original training set with only 2019 at 90 GHz from PA5 data. The test set remained the same throughout, in order to ensure that the accuracy was not increasing due to removing difficult-to-classify test points.
@@
The results from this can be seen in Figure 10 where the left plot shows the confusion matrix that arises from classifications based on the initial training set. The middle and right plots show the confusion matrices where additional training points have been chosen though active learning (middle) or random selection (right). In both cases, the overall accuracy of classification increases modestly, by and percentage points, respectively. We note that the classification accuracy for the PS category increased significantly, by close to 14 percentage points, when using an active learning adjustment and by only 5 percentage points when supplementing the training set randomly.
@@
Similarly, the accuracy for PS+ increased by 11 percentage points when utilizing active learning, as opposed to 5 percentage points with random selection. This is promising: as the volume of data increases with successive observing seasons, new detector arrays, and looking to future telescopes such as SO, we are now motivated to use active learning to decrease the amount of manual labeling needed in a training set. Active learning also allows us to ensure that we are choosing test points that are difficult for our algorithm to classify, and therefore gaining a better understanding of its performance.

### v2

@@ 4.2.2 Normalized Maximum
To derive this statistic, we compute the histogram of the distribution of the detectors in the direction of the focal plane for a given glitch. The histogram bin with the largest number of detectors and the adjacent bins are added together and then normalized by the total number of detectors present in the glitch. We compute the histogram of detectors with 10 bins, resulting in the width of an individual bin in the histogram being small when the detectors are clustered and larger if there is spread in positions. Figure 3 shows examples for both the PS and PS+ categories, where the detectors being considered are shown as coral compared to all the detectors in the snippet shown in purple. Note that in both examples, there is only one bin beside the maximum bin, thus only two bins are used for the normalized maximum statistic. For the PS+, this is because the maximum bin is the first bin, whereas for the PS it is due to the detectors that would have been in the bin to the right of the maximum being cut during the detectors cuts which happen before the glitch cuts. The PS object, has only 27% of the detectors within the maximum and adjacent bins, compared to 59% of detectors within the maximum and adjacent bins for the PS+ object.
@@
In general, we expect the spread of affected detectors in the direction to fall within 0.1, or a few beams, for an astrophysical source. This is not true for a general electronic glitch or cosmic ray, which makes this metric useful for identifying astrophysical sources. We compute the fraction of detectors in the direction of the focal plane that are within 0.1 of the histogram bin that contains the maximum number of detectors. Figure 3 shows examples for both the PS and PS+ categories for both the normalized maximum and normalized maximum within 0.1 to compare these statistics.
Similar to the previously defined statistic, we expect a small range of detectors to observe a point source as the telescope scans over the source. Therefore, the number of detectors within 0.1 of the position of the maximum histogram bin should be close to the total number of detectors in the snippet (and therefore our statistic normalized over the total number of detectors should be very close to unity, as shown in Figures 2 and 3). This makes the statistic very useful for determining if a glitch is a point source. A cosmic ray will hit a small cluster of detectors, resulting in a small spread in the direction. We still expect the fraction of the total number of detectors within 0.1 of the position to be close to unity, but there can be deviations from this if multiple cosmic rays hit the focal plane at once. The expected distribution of detectors varies quite significantly depending on the type of electronic glitch is occurring, but we find that the statistic is generally smaller for electronic glitches than for the PS class. For a PS+ object, the value of this statistic will differ from what is seen for an isolated source depending once again on where the additional glitch occurs on the focal plane relative to the source location: if the glitch occurs near the source, we similarly expect many detectors to be included, leading to a large fraction of the detectors to be within 0.1 of the maximum. Conversely, if the glitch occurs far away on the focal plane, this will result in a smaller fraction of detectors within 0.1 of the maximum and a smaller value for our statistic. This statistical metric is therefore useful in identifying PS+ objects where the coincident glitches are located elsewhere in the detector array and thus would be able to be easily separated in the snippets from the point source and for further analysis of the source.
@@ 4.3 Correlation
In order to determine how the incident flux is correlated across detectors (which can help distinguish between the PS and CR classes), we compute the Pearson correlation coefficient of each detector pair ordered by their position across the focal plane. We then use the mean of the absolute value of these correlation coefficients from all detector pairs in the snippet as a feature for our random forest. Given that detectors observe the point source sequentially as the telescope scans across the sky, we only expect a large value of this correlation coefficient between detectors that are either close to or beside the main detectors on the focal plane that observe the source. This scenario leads to a block diagonal correlation matrix of high values across the array when the detectors are ordered by their position (see Figure 4 for an illustration) but a low overall degree of correlation across all detectors. For cosmic rays, we expect a high level of correlation across the array, because the detectors are all affected by the cosmic ray at the same time, shown in Figure 4. For electronic glitches that cause a strong signal that affects multiple detectors simultaneously, we expect a large value of the correlation coefficient during, e.g., a large-area heat-up. Conversely, if the glitch is similar to a random noise fluctuation, we would expect a low value of the correlation coefficient. For PS+ objects, if the additional glitch is a cosmic ray, we expect a slightly higher correlation than would be observed for a point source, as the cosmic ray portion of the TODs for each detector will overlap while affected by the cosmic ray. If the additional glitch is of electronic or other origin, the effect varies.
@@ 4.4 Time-lag
While the absolute value of the correlation described above indicates the spatial correlation between detectors that see a given glitch, the temporal correlation (or time-lag) is also a key statistic that we use in classification. The time-lag between detectors is the shift in time that maximizes the cross-correlation between the two detector TODs over the snippet. For two detectors and , it is defined following 35:
where “DFT” is the discrete Fourier transform, the overbar denotes the complex conjugate, and is the timestream of the detector as a function of time. We take the absolute value, since the sign of the shift (backwards or forwards) does not contain information relevant to the type of the signal. We compute this time-lag for all detector pairs in the snippet, and then use the mean of the absolute values to summarize the glitch.
For a point source, each detector observes the source sequentially as the telescope scans across the sky. We therefore expect a time-lag that is equal to the separation in azimuth of the two detectors divided by the scanning speed. In practice, we can get different values of the time-lag based on how noisy the TOD is, especially for low amplitude sources. This can be used in future work to differentiate an astrophysical point source from a moving source, such as a satellite. Conversely, we expect a time-lag close to zero for cosmic rays, since the cosmic ray timescale is fast compared to the readout rate, which means that the digitized signatures reside essentially on top of each other in the detector timestream. This can be seen in Figure 5, where a time-lag matrix for detector pairs, ordered by their position, is shown for a cosmic ray on the top and a point source on the bottom. As expected, the time-lag is zero for almost every detector pair for the cosmic ray. For the EG category, it will cause varied time-lags depending on the nature of the glitch. Generally, something such as a focal plane heating event or another electronic glitch that affects all the detectors at the same time will result in a zero time-lag as the detectors experience the effect at the same time. Finally, for the PS+ class, if the other glitch is a cosmic ray, we expect a slightly smaller time-lag than just a point source as part of the TODs will be overlapping with a zero time-lag. If the other glitch is of an electronic or other origin the effect varies, but generally speaking it increases the time-lag, as we show in Figure 2.
@@ 4.5 Number of Peaks
In a given snippet, some glitches, in particular glitches from PS objects and CR objects, will appear as “peaks” in the TOD. Thus, the number of signal peaks at unique times is used as one of our summary statistics, as it is particularly useful in discerning a cosmic ray from an electronic glitch. In order to compute the number of peaks in a snippet, we first smooth the TODs for the snippet using a boxcar smoothing kernel of width three time samples to remove small fluctuations due to random noise. We then combine the TODs from each detector into a single timestream , as:
@@
5 Training and Testing the Random Forest
@@ 5.1 Labeling of Objects in the Training and Test Sets
In order to create a large enough training set for the random forest classifier, we needed an efficient method to facilitate manual labeling of the snippets, via a graphical user interface (GUI).
We use the Python-driven platform developed by the Zooniverse collaboration222www.zooniverse.org to create the labeling GUI. This Zooniverse platform includes a well-established data pipeline and a user-friendly interface, as can be seen in Figure 7. The platform we employ was designed specifically for use by members of the ACT collaboration, to assist in both training and verification of the random forest.
To facilitate classifications by non-expert users, we created a user tutorial specific to ACT glitch data, with examples related to the key features of each type of glitch and notes on pitfalls in classifying glitches.333We separately included instructions on setting up a Zooniverse project from scratch, see online documentation here.
@@ 5.2 Forming the Training and Test Sets
Once we have obtained a representative sample for training and testing the random forest, we need to ensure we are robust to changes in the snippets from different years (observation seasons), observational frequency bands, detector array configurations, etc., which requires further refinement of our test/training split. In order to ensure we are robust to these changes, we use an active learning (AL) prescription, in particular the modAL 18 learning framework. AL is a machine learning approach where unlabeled objects are selected for manual classification/labeling depending on specific criteria of the object (e.g. the classification ambiguity between types in a given set, or an estimate of the Bayesian classification of the object type), as opposed to selecting the point for labeling randomly from the unlabeled set. We employed the modAL package to choose objects using a selection criterion based on which objects are most ambiguous, or in other words, objects for which the maximum classification probabilities are the lowest. In doing so, we can leverage human/manual classifications where our random forest classifier might be most uncertain. This criterion for selecting objects for manual classification is known as “uncertainty sampling”. Figure 8 shows a diagram of the process.
@@
From the initial random forest classification, each glitch has four probabilities related to the four classification categories: PS, PS+, CR, and EG. The most likely (highest) probabilities for all glitches are ranked, and the uncertainty sampler takes the lowest value of those probabilities. In other words, the uncertainty sampler selects a glitch that the random forest classifier is the least confident in (or most uncertain about) to manually label and therefore add to the AL training set 30. This is shown in the bottom right-hand section of Figure 8.
@@
To demonstrate the effects of AL, we tested the random forest on a set of 851 labeled glitches from 2021 at 90 GHz and 150 GHz for detector arrays PA4 and PA5. We begin with a training set composed of 2280 glitches from 2019 at 90 GHz in PA5. We added 300 training points using two methods, AL and then random selection, that were chosen from a set of 2555 glitches from 2021 at 90 GHz and 150 GHz in PA4 and PA5. Note that the 851 and 2555 glitches were chosen by a 25/75 % random split from an initial data set of 3406 glitches. We chose to add 300 samples as the accuracy increase for the PS and PS+ categories, which we care most about, plateau. Additionally, we want to test realistic settings for the future where ideally we will need to label a small number of additional training points for new seasons of data, additional arrays, etc. We trained three random forests using the three training sets: original, addition with AL, and random addition. The test set is then classified using each of the forests to compare their performance. It is important to note that the test set and additional training points are comprised of data from different arrays, frequencies, and years which could have different noise properties or detector performance than the original training set with only data from 2019 at 90 GHz for PA5. The test set remained the same throughout, in order to ensure that the accuracy was not increasing due to removing difficult-to-classify test points. The class distribution for the test, original training, training points that were available to add, training with AL addition, and training with random addition sets can be seen in Table 1.444All of the training and test sets can be found at https://phy-act1.princeton.edu/public/data/dr6_tod_v1/.
@@
The results can be seen in Figure 9 where the left plot shows the confusion matrix that arises from classifications based on the initial training set. Each value in the confusion matrix indicates the decimal percentage of each glitch category on the -axis being predicted as the category shown on the -axis. For perfect performance, the resulting confusion matrix would be an identity matrix. The middle and right plots show the confusion matrices where additional training points have been chosen through AL (middle) or random selection (right). In both cases, the overall accuracy of classification increases modestly, by and points, respectively. We note that the classification accuracy for the PS category increased significantly, by close to 14% points, when using an AL adjustment and by only 5% points when supplementing the training set randomly.
@@
Similarly, the accuracy for PS+ increased by 11% points when utilizing AL, as opposed to 2% points with random selection. This is promising: As the volume of data increases with successive observing seasons, new detector arrays, and looking to future telescopes such as SO, we are now motivated to use AL to decrease the amount of manual labeling needed in a training set. AL also allows us to ensure that we are choosing test points that are difficult for the trained random forest to classify, and therefore gaining a better understanding of its performance.
5.3 Classification Performance Metrics
| Number of Snippets | |||||
| Data Set | Total | PS | PS+ | CR | EG |
| Test | 851 | 63 | 44 | 363 | 381 |
| Original training | 2280 | 239 | 121 | 1040 | 880 |
| Available for addition | 2555 | 170 | 156 | 1068 | 1161 |
| AL training | 2580 | 272 | 167 | 1119 | 1022 |
| Random training | 2580 | 255 | 138 | 1142 | 1045 |
To test the performance of the classifier, we also examined the precision, recall, and F-1 score for each of the classification categories. As illustrated in Equation 4, precision is the proportion of glitches predicted to be within a category that were predicted correctly, while recall measures how many glitches of a certain category were correctly predicted. A high precision score therefore indicates a low false positive rate (few objects incorrectly identified), whereas a high recall indicates a low false negative rate (few objects missed). The F-1 score is a harmonic mean of the precision and recall of the forest. They are given by:
where “tp” is the number of true positives, “fp” is the number of false positives, “fn” is the number of false negatives, is the precision, is the recall, and “F-1” is the F-1 score.
5.4 Fine-tuning the Random Forest
We explored the performance of the random forest as a function of the number of trees and the “max depth” (which is the maximum number of times each decision tree is allowed to split) of the forest classifier. The number of decision trees used was varied between 25 and 175 (preliminary tests outside that range showed minimal impact in performance) and the max depth varied within an initial range of 2 to 32. The performance was assessed using the overall accuracy of the classifier, the overall out-of-bag (OOB) score, and F-1 score. The random forest approach uses bootstrap aggregation to train each tree, a process whereby the training set for each decision tree is produced by sampling from the original training set with replacement. This avoids overfitting of the random forest to the training data, which can lead to a model that is does not perform well on new (test) data. The accuracy of each tree is computed with the samples that were not included in its bootstrapped training set. The average accuracy score from all the trees is known as the OOB score and is used to characterize the ability of a method like random forests to make predictions on data that were not included in the training set 24. The performance metrics discussed above plateau and do not improve when the forest reaches a size of 50 trees and a max depth value of 15; therefore, we fixed those hyperparameters of the random forest method to these values.
5.5 Feature Importance
When using the various statistics described in Section 4, it is important to determine which of the features are more instructive in the classification step, to ensure that we understand how the classifier weights them. Low-importance features can potentially be removed from a feature set, and high-importance features can be examined to identify the physics behind why the features lead to more accurate classifications. It also aids in our understanding of misclassified glitches; see Section 7.1 for more details.